# REAP the Experts: Why Pruning Prevails for One-Shot MoE compression

### v1

@@ Appendix E Additional results
ERNIE-4.5-21B-
A3B-PT
Baseline 0.902 0.866 0.910 0.765 0.861 0.231 0.546
25% Merging M-SMoE 0.774 $\pm$ 0.011 0.730 $\pm$ 0.009 0.768 $\pm$ 0.015 0.647 $\pm$ 0.017 0.730 $\pm$ 0.005 0.194 $\pm$ 0.022 0.462 $\pm$ 0.011
HC-SMoE 0.837 $\pm$ 0.007 0.805 $\pm$ 0.000 0.827 $\pm$ 0.003 0.696 $\pm$ 0.008 0.791 $\pm$ 0.004 0.207 $\pm$ 0.008 0.499 $\pm$ 0.003
Pruning Frequency 0.890 $\pm$ 0.006 0.846 $\pm$ 0.009 0.837 $\pm$ 0.010 0.709 $\pm$ 0.010 0.820 $\pm$ 0.006 0.151 $\pm$ 0.096 0.486 $\pm$ 0.045
EAN 0.890 $\pm$ 0.006 0.848 $\pm$ 0.011 0.840 $\pm$ 0.006 0.727 $\pm$ 0.004 0.826 $\pm$ 0.004 0.161 $\pm$ 0.111 0.494 $\pm$ 0.054
REAP 0.892 $\pm$ 0.009 0.854 $\pm$ 0.012 0.876 $\pm$ 0.000 0.738 $\pm$ 0.003 0.840 $\pm$ 0.005 0.167 $\pm$ 0.124 0.504 $\pm$ 0.060
50% Merging M-SMoE 0.104 $\pm$ 0.022 0.100 $\pm$ 0.029 0.239 $\pm$ 0.036 0.207 $\pm$ 0.040 0.162 $\pm$ 0.012 0.024 $\pm$ 0.008 0.093 $\pm$ 0.008
HC-SMoE 0.425 $\pm$ 0.004 0.404 $\pm$ 0.007 0.608 $\pm$ 0.018 0.511 $\pm$ 0.011 0.487 $\pm$ 0.008 0.082 $\pm$ 0.015 0.285 $\pm$ 0.009
Pruning Frequency 0.699 $\pm$ 0.031 0.640 $\pm$ 0.022 0.696 $\pm$ 0.014 0.584 $\pm$ 0.006 0.655 $\pm$ 0.015 0.083 $\pm$ 0.066 0.369 $\pm$ 0.025
EAN 0.675 $\pm$ 0.019 0.642 $\pm$ 0.009 0.713 $\pm$ 0.015 0.591 $\pm$ 0.016 0.655 $\pm$ 0.014 0.112 $\pm$ 0.064 0.384 $\pm$ 0.035
REAP 0.797 $\pm$ 0.009 0.764 $\pm$ 0.007 0.767 $\pm$ 0.017 0.644 $\pm$ 0.013 0.743 $\pm$ 0.008 0.137 $\pm$ 0.119 0.440 $\pm$ 0.064
Qwen3-30B-A3B Baseline 0.927 0.884 0.881 0.743 0.859 0.302 0.581
25% Merging M-SMoE 0.878 $\pm$ 0.012 0.833 $\pm$ 0.007 0.849 $\pm$ 0.007 0.728 $\pm$ 0.007 0.822 $\pm$ 0.004 0.293 $\pm$ 0.017 0.558 $\pm$ 0.006
HC-SMoE 0.866 $\pm$ 0.011 0.805 $\pm$ 0.016 0.832 $\pm$ 0.006 0.698 $\pm$ 0.005 0.800 $\pm$ 0.004 0.258 $\pm$ 0.000 0.529 $\pm$ 0.002
Pruning Frequency 0.921 $\pm$ 0.006 0.874 $\pm$ 0.007 0.868 $\pm$ 0.000 0.735 $\pm$ 0.003 0.849 $\pm$ 0.004 0.302 $\pm$ 0.011 0.576 $\pm$ 0.004
EAN 0.909 $\pm$ 0.006 0.864 $\pm$ 0.004 0.859 $\pm$ 0.009 0.729 $\pm$ 0.008 0.840 $\pm$ 0.004 0.311 $\pm$ 0.018 0.576 $\pm$ 0.010
REAP 0.917 $\pm$ 0.007 0.876 $\pm$ 0.004 0.853 $\pm$ 0.002 0.727 $\pm$ 0.006 0.843 $\pm$ 0.002 0.308 $\pm$ 0.015 0.575 $\pm$ 0.008
50% Merging M-SMoE 0.687 $\pm$ 0.013 0.638 $\pm$ 0.004 0.618 $\pm$ 0.004 0.541 $\pm$ 0.007 0.621 $\pm$ 0.006 0.205 $\pm$ 0.019 0.413 $\pm$ 0.007
HC-SMoE 0.577 $\pm$ 0.023 0.541 $\pm$ 0.013 0.631 $\pm$ 0.010 0.546 $\pm$ 0.004 0.574 $\pm$ 0.010 0.185 $\pm$ 0.018 0.379 $\pm$ 0.005
Pruning Frequency 0.787 $\pm$ 0.016 0.756 $\pm$ 0.022 0.692 $\pm$ 0.016 0.579 $\pm$ 0.016 0.704 $\pm$ 0.017 0.236 $\pm$ 0.025 0.470 $\pm$ 0.021
EAN 0.886 $\pm$ 0.025 0.837 $\pm$ 0.020 0.798 $\pm$ 0.006 0.669 $\pm$ 0.008 0.798 $\pm$ 0.013 0.306 $\pm$ 0.003 0.552 $\pm$ 0.005
REAP 0.919 $\pm$ 0.007 0.870 $\pm$ 0.004 0.805 $\pm$ 0.009 0.692 $\pm$ 0.008 0.821 $\pm$ 0.003 0.293 $\pm$ 0.003 0.557 $\pm$ 0.001
Mixtral-8x7B-
Instruct-v0.1
Baseline 0.524 0.476 0.556 0.463 0.505 0.123 0.314
25% Merging M-SMoE 0.315 $\pm$ 0.007 0.270 $\pm$ 0.015 0.446 $\pm$ 0.007 0.380 $\pm$ 0.015 0.353 $\pm$ 0.008 0.033 $\pm$ 0.010 0.193 $\pm$ 0.008
HC-SMoE 0.439 $\pm$ 0.028 0.386 $\pm$ 0.020 0.530 $\pm$ 0.022 0.441 $\pm$ 0.007 0.449 $\pm$ 0.005 0.110 $\pm$ 0.010 0.279 $\pm$ 0.002
Pruning Frequency 0.400 $\pm$ 0.034 0.358 $\pm$ 0.035 0.541 $\pm$ 0.006 0.453 $\pm$ 0.012 0.438 $\pm$ 0.018 0.099 $\pm$ 0.014 0.269 $\pm$ 0.004
EAN 0.413 $\pm$ 0.027 0.366 $\pm$ 0.024 0.477 $\pm$ 0.009 0.409 $\pm$ 0.013 0.416 $\pm$ 0.015 0.111 $\pm$ 0.006 0.264 $\pm$ 0.006
REAP 0.439 $\pm$ 0.018 0.370 $\pm$ 0.007 0.535 $\pm$ 0.011 0.452 $\pm$ 0.011 0.449 $\pm$ 0.002 0.102 $\pm$ 0.010 0.275 $\pm$ 0.005
50% Merging M-SMoE 0.085 $\pm$ 0.026 0.076 $\pm$ 0.022 0.139 $\pm$ 0.121 0.118 $\pm$ 0.102 0.091 $\pm$ 0.079 0.004 $\pm$ 0.006 0.047 $\pm$ 0.037
HC-SMoE 0.175 $\pm$ 0.015 0.146 $\pm$ 0.000 0.335 $\pm$ 0.026 0.282 $\pm$ 0.031 0.235 $\pm$ 0.018 0.013 $\pm$ 0.008 0.124 $\pm$ 0.008
Pruning Frequency 0.187 $\pm$ 0.015 0.148 $\pm$ 0.007 0.342 $\pm$ 0.016 0.287 $\pm$ 0.012 0.241 $\pm$ 0.007 0.023 $\pm$ 0.004 0.132 $\pm$ 0.003
EAN 0.220 $\pm$ 0.006 0.189 $\pm$ 0.006 0.375 $\pm$ 0.020 0.325 $\pm$ 0.015 0.277 $\pm$ 0.005 0.031 $\pm$ 0.011 0.154 $\pm$ 0.007
REAP 0.232 $\pm$ 0.018 0.193 $\pm$ 0.013 0.274 $\pm$ 0.106 0.241 $\pm$ 0.087 0.235 $\pm$ 0.056 0.035 $\pm$ 0.003 0.135 $\pm$ 0.027
Llama-4-Scout-
17B-16E-
Instruct
Baseline 0.829 0.768 0.788 0.640 0.757 0.341 0.549
25% Merging M-SMoE 0.823 0.762 0.786 0.635 0.752 0.324 0.538
HC-SMoE 0.787 0.738 0.735 0.587 0.712 0.148 0.430
Pruning Frequency 0.835 0.768 0.788 0.630 0.755 0.317 0.536
EAN 0.823 0.762 0.804 0.648 0.759 0.328 0.544
REAP 0.829 0.787 0.788 0.622 0.756 0.242 0.499
50% Merging M-SMoE 0.787 0.732 0.762 0.614 0.723 0.187 0.455
HC-SMoE 0.604 0.530 0.500 0.399 0.508 0.077 0.293
Pruning Frequency 0.823 0.756 0.751 0.595 0.731 0.223 0.477
EAN 0.805 0.744 0.754 0.601 0.726 0.209 0.468
REAP 0.841 0.768 0.762 0.624 0.749 0.248 0.499
GLM-4.5-Air Baseline 0.848 0.829 0.860 0.743 0.820 0.374 0.597
25% Merging M-SMoE 0.866 0.793 0.807 0.659 0.781 0.330 0.555
HC-SMoE 0.872 0.805 0.825 0.669 0.793 0.363 0.578
Pruning Frequency 0.848 0.811 0.854 0.706 0.805 0.341 0.573
EAN 0.872 0.817 0.876 0.720 0.821 0.374 0.597
REAP 0.866 0.805 0.828 0.677 0.794 0.390 0.592
50% Merging M-SMoE 0.518 0.500 0.519 0.437 0.493 0.099 0.296
HC-SMoE 0.707 0.659 0.706 0.577 0.662 0.220 0.441
Pruning Frequency 0.628 0.573 0.534 0.450 0.546 0.104 0.325
EAN 0.841 0.780 0.807 0.661 0.773 0.253 0.513
REAP 0.878 0.841 0.712 0.587 0.755 0.352 0.553
Qwen3-Coder-
480B-A35B-
Instruct-FP8
Baseline 0.951 0.890 0.923 0.791 0.889 0.431 $\pm$ 0.011 0.660
25% Pruning Frequency 0.884 0.805 0.810 0.669 0.792 0.296 $\pm$ 0.017 0.544
EAN 0.939 0.878 0.911 0.775 0.876 0.419 $\pm$ 0.015 0.647
REAP 0.957 0.890 0.917 0.772 0.884 0.416 $\pm$ 0.013 0.650
50% Pruning Frequency 0.020 0.012 0.007 0.003 0.011 0.012 $\pm$ 0.001 0.011
EAN 0.915 0.841 0.854 0.714 0.831 0.382 $\pm$ 0.012 0.607
REAP 0.939 0.872 0.910 0.772 0.873 0.415 $\pm$ 0.015 0.644
Kimi-K2-
Instruct-
W4A16
Baseline 0.963 0.921 0.913 0.735 0.883 0.434 0.659
25% Pruning Frequency 0.530 0.463 0.595 0.508 0.524 0.082 0.303
EAN 0.909 0.860 0.857 0.698 0.831 0.379 0.605
REAP 0.957 0.921 0.918 0.759 0.889 0.440 0.664
50% Pruning Frequency 0.098 0.079 0.175 0.146 0.124 0.000 0.062
EAN 0.866 0.811 0.780 0.632 0.772 0.253 0.513
REAP 0.915 0.884 0.899 0.754 0.863 0.429 0.646
@@
ERNIE-4.5-21B-
A3B-PT
Baseline 0.861 0.231 0.546 0.564 0.782 0.873 0.813 0.737 0.462 0.812 0.724 0.721
25% Merging M-SMoE 0.065 0.016 0.041 0.497 0.729 0.860 0.723 0.602 0.424 0.801 0.699 0.667
HC-SMoE 0.403 0.099 0.251 0.515 0.728 0.860 0.745 0.649 0.428 0.794 0.694 0.677
Pruning Frequency 0.274 0.000 0.137 0.515 0.735 0.841 0.719 0.588 0.382 0.791 0.683 0.657
EAN 0.282 0.000 0.141 0.528 0.750 0.853 0.790 0.558 0.442 0.783 0.706 0.676
REAP 0.242 0.023 0.133 0.490 0.716 0.855 0.783 0.656 0.452 0.809 0.723 0.685
@@
REAP 0.033 0.000 0.016 0.406 0.612 0.754 0.654 0.468 0.396 0.718 0.656 0.583
Qwen3-30B-A3B Baseline 0.859 0.302 0.581 0.563 0.790 0.887 0.778 0.779 0.454 0.816 0.702 0.721
@@
REAP 0.006 0.000 0.003 0.421 0.640 0.837 0.653 0.495 0.388 0.704 0.635 0.596
Mixtral-8x7B-
Instruct-v0.1
Baseline 0.505 0.123 0.314 0.650 0.842 0.887 0.861 0.691 0.496 0.722 0.740 0.736
25% Merging M-SMoE 0.320 0.044 0.182 0.532 0.775 0.828 0.746 0.529 0.424 0.603 0.632 0.634
HC-SMoE 0.420 0.121 0.271 0.608 0.811 0.876 0.838 0.631 0.484 0.736 0.726 0.714
Pruning Frequency 0.396 0.070 0.233 0.612 0.816 0.868 0.836 0.593 0.482 0.675 0.739 0.703
EAN 0.399 0.092 0.246 0.613 0.814 0.875 0.842 0.613 0.498 0.690 0.733 0.710
REAP 0.415 0.077 0.246 0.606 0.807 0.875 0.835 0.633 0.486 0.791 0.709 0.718
@@
HC-SMoE 0.174 0.033 0.103 0.540 0.764 0.862 0.795 0.544 0.448 0.675 0.709 0.667
Pruning Frequency 0.173 0.008 0.090 0.504 0.739 0.793 0.771 0.463 0.426 0.675 0.646 0.627
EAN 0.139 0.008 0.074 0.550 0.756 0.842 0.804 0.529 0.460 0.726 0.716 0.673
REAP 0.167 0.012 0.089 0.525 0.774 0.856 0.794 0.533 0.454 0.751 0.688 0.672
@@
Figure˜A5 plots non-agentic coding and MC accuracy versus compressed model size. Figure˜6(a) depict the proportion of singleton clusters for HC-SMoE and M-SMoE. Figure˜6(b) plots accuracy vs. maximum cluster sizes when the maximum cardinality of clusters is restricted. Figures˜A7 and A8 show the importance of using domain-specific calibration data, particularly at high compression ratios.
Table˜A7 presents the complete $\tau^{2}$-bench results across three domains (Retail, Airline, and Telecom) for the baseline model and REAP compression at 25% and 50% levels. The results show passˆk metrics for k=1, 2, and 3, demonstrating the impact of pruning on evaluating conversational agents, specifically designed to test their ability to collaborate with a user in real-world scenarios.
@@
Figure A6: (6(a)) Average proportion of singleton clusters vs. model for HC-SMoE and M-SMoE. We find that the clustering algorithms used by our baseline merging methods tend to generate a high proportion of singleton clusters containing just a single expert. In order to achieve the desired compression ratio, the large number of singletons conversely results in some clusters which contain many experts, in some cases $N/2+1$ experts for a layer with $N$ experts are grouped into a single cluster. (6(b)) Accuracy vs. maximum cluster size using M-SMoE to compress 50% of experts in Qwen3-30B. While MC accuracy remains stable up to a maximum cluster size of 4, generative coding capabilities are severely diminished by restricting the clustering algorithm.

### v2

@@ Appendix F Additional results
ERNIE-4.5-21B- A3B-PT Baseline 0.902 0.866 0.910 0.765 0.815 0.231 0.523
25% Merging M-SMoE 0.774 $\pm$ 0.011 0.730 $\pm$ 0.009 0.768 $\pm$ 0.015 0.647 $\pm$ 0.017 0.688 $\pm$ 0.007 0.194 $\pm$ 0.022 0.441 $\pm$ 0.011
HC-SMoE 0.837 $\pm$ 0.007 0.805 $\pm$ 0.000 0.827 $\pm$ 0.003 0.696 $\pm$ 0.008 0.750 $\pm$ 0.004 0.207 $\pm$ 0.008 0.479 $\pm$ 0.003
Pruning Frequency 0.890 $\pm$ 0.006 0.846 $\pm$ 0.009 0.837 $\pm$ 0.010 0.709 $\pm$ 0.010 0.777 $\pm$ 0.009 0.151 $\pm$ 0.096 0.464 $\pm$ 0.044
EAN 0.890 $\pm$ 0.006 0.848 $\pm$ 0.011 0.840 $\pm$ 0.006 0.727 $\pm$ 0.004 0.787 $\pm$ 0.007 0.161 $\pm$ 0.111 0.474 $\pm$ 0.053
REAP 0.909 $\pm$ 0.012 0.868 $\pm$ 0.004 0.866 $\pm$ 0.006 0.735 $\pm$ 0.002 0.801 $\pm$ 0.002 0.223 $\pm$ 0.008 0.512 $\pm$ 0.005
50% Merging M-SMoE 0.104 $\pm$ 0.022 0.100 $\pm$ 0.029 0.239 $\pm$ 0.036 0.207 $\pm$ 0.040 0.153 $\pm$ 0.015 0.024 $\pm$ 0.008 0.089 $\pm$ 0.009
HC-SMoE 0.425 $\pm$ 0.004 0.404 $\pm$ 0.007 0.608 $\pm$ 0.018 0.511 $\pm$ 0.011 0.458 $\pm$ 0.008 0.082 $\pm$ 0.015 0.270 $\pm$ 0.008
Pruning Frequency 0.699 $\pm$ 0.031 0.640 $\pm$ 0.022 0.696 $\pm$ 0.014 0.584 $\pm$ 0.006 0.612 $\pm$ 0.014 0.083 $\pm$ 0.066 0.348 $\pm$ 0.026
EAN 0.675 $\pm$ 0.019 0.642 $\pm$ 0.009 0.713 $\pm$ 0.015 0.591 $\pm$ 0.016 0.617 $\pm$ 0.012 0.112 $\pm$ 0.064 0.364 $\pm$ 0.034
REAP 0.787 $\pm$ 0.038 0.752 $\pm$ 0.035 0.749 $\pm$ 0.005 0.638 $\pm$ 0.013 0.695 $\pm$ 0.024 0.187 $\pm$ 0.005 0.441 $\pm$ 0.014
Qwen3-30B-A3B Baseline 0.927 0.884 0.881 0.743 0.814 0.302 0.558
25% Merging M-SMoE 0.878 $\pm$ 0.012 0.833 $\pm$ 0.007 0.849 $\pm$ 0.007 0.728 $\pm$ 0.007 0.781 $\pm$ 0.007 0.293 $\pm$ 0.017 0.537 $\pm$ 0.006
HC-SMoE 0.866 $\pm$ 0.011 0.805 $\pm$ 0.016 0.832 $\pm$ 0.006 0.698 $\pm$ 0.005 0.752 $\pm$ 0.006 0.258 $\pm$ 0.000 0.505 $\pm$ 0.003
Pruning Frequency 0.921 $\pm$ 0.006 0.874 $\pm$ 0.007 0.868 $\pm$ 0.000 0.735 $\pm$ 0.003 0.805 $\pm$ 0.005 0.302 $\pm$ 0.011 0.553 $\pm$ 0.003
EAN 0.909 $\pm$ 0.006 0.864 $\pm$ 0.004 0.859 $\pm$ 0.009 0.729 $\pm$ 0.008 0.797 $\pm$ 0.005 0.311 $\pm$ 0.018 0.554 $\pm$ 0.010
REAP 0.911 $\pm$ 0.004 0.870 $\pm$ 0.004 0.847 $\pm$ 0.004 0.725 $\pm$ 0.008 0.797 $\pm$ 0.004 0.304 $\pm$ 0.003 0.551 $\pm$ 0.004
50% Merging M-SMoE 0.687 $\pm$ 0.013 0.638 $\pm$ 0.004 0.618 $\pm$ 0.004 0.541 $\pm$ 0.007 0.590 $\pm$ 0.005 0.205 $\pm$ 0.019 0.397 $\pm$ 0.007
HC-SMoE 0.577 $\pm$ 0.023 0.541 $\pm$ 0.013 0.631 $\pm$ 0.010 0.546 $\pm$ 0.004 0.543 $\pm$ 0.005 0.185 $\pm$ 0.018 0.364 $\pm$ 0.007
Pruning Frequency 0.787 $\pm$ 0.016 0.756 $\pm$ 0.022 0.692 $\pm$ 0.016 0.579 $\pm$ 0.016 0.668 $\pm$ 0.019 0.236 $\pm$ 0.025 0.452 $\pm$ 0.022
EAN 0.886 $\pm$ 0.025 0.837 $\pm$ 0.020 0.798 $\pm$ 0.006 0.669 $\pm$ 0.008 0.753 $\pm$ 0.011 0.306 $\pm$ 0.003 0.530 $\pm$ 0.004
REAP 0.917 $\pm$ 0.013 0.858 $\pm$ 0.015 0.818 $\pm$ 0.008 0.703 $\pm$ 0.004 0.780 $\pm$ 0.006 0.302 $\pm$ 0.000 0.541 $\pm$ 0.003
Mixtral-8x7B- Instruct-v0.1 Baseline 0.524 0.476 0.556 0.463 0.469 0.123 0.296
25% Merging M-SMoE 0.315 $\pm$ 0.007 0.270 $\pm$ 0.015 0.446 $\pm$ 0.007 0.380 $\pm$ 0.015 0.325 $\pm$ 0.015 0.033 $\pm$ 0.010 0.179 $\pm$ 0.011
HC-SMoE 0.439 $\pm$ 0.028 0.386 $\pm$ 0.020 0.530 $\pm$ 0.022 0.441 $\pm$ 0.007 0.414 $\pm$ 0.007 0.110 $\pm$ 0.010 0.262 $\pm$ 0.001
Pruning Frequency 0.400 $\pm$ 0.034 0.358 $\pm$ 0.035 0.541 $\pm$ 0.006 0.453 $\pm$ 0.012 0.405 $\pm$ 0.019 0.099 $\pm$ 0.014 0.252 $\pm$ 0.004
EAN 0.413 $\pm$ 0.027 0.366 $\pm$ 0.024 0.477 $\pm$ 0.009 0.409 $\pm$ 0.013 0.388 $\pm$ 0.015 0.111 $\pm$ 0.006 0.249 $\pm$ 0.006
REAP 0.445 $\pm$ 0.016 0.388 $\pm$ 0.025 0.548 $\pm$ 0.010 0.470 $\pm$ 0.011 0.429 $\pm$ 0.008 0.097 $\pm$ 0.006 0.263 $\pm$ 0.007
50% Merging M-SMoE 0.085 $\pm$ 0.026 0.076 $\pm$ 0.022 0.139 $\pm$ 0.121 0.118 $\pm$ 0.102 0.127 $\pm$ 0.011 0.004 $\pm$ 0.006 0.063 $\pm$ 0.005
HC-SMoE 0.175 $\pm$ 0.015 0.146 $\pm$ 0.000 0.335 $\pm$ 0.026 0.282 $\pm$ 0.031 0.214 $\pm$ 0.015 0.013 $\pm$ 0.008 0.114 $\pm$ 0.007
Pruning Frequency 0.187 $\pm$ 0.015 0.148 $\pm$ 0.007 0.342 $\pm$ 0.016 0.287 $\pm$ 0.012 0.217 $\pm$ 0.007 0.023 $\pm$ 0.004 0.120 $\pm$ 0.004
EAN 0.220 $\pm$ 0.006 0.189 $\pm$ 0.006 0.375 $\pm$ 0.020 0.325 $\pm$ 0.015 0.257 $\pm$ 0.005 0.031 $\pm$ 0.011 0.144 $\pm$ 0.006
REAP 0.258 $\pm$ 0.019 0.220 $\pm$ 0.016 0.381 $\pm$ 0.003 0.331 $\pm$ 0.008 0.275 $\pm$ 0.011 0.055 $\pm$ 0.010 0.165 $\pm$ 0.001
Llama-4-Scout- 17B-16E- Instruct Baseline 0.829 0.768 0.788 0.640 0.704 0.341 0.522
25% Merging M-SMoE 0.823 0.762 0.786 0.635 0.699 0.324 0.511
HC-SMoE 0.787 0.738 0.735 0.587 0.663 0.148 0.405
Pruning Frequency 0.835 0.768 0.788 0.630 0.699 0.317 0.508
EAN 0.823 0.762 0.804 0.648 0.705 0.328 0.517
REAP 0.829 0.787 0.788 0.622 0.704 0.242 0.473
50% Merging M-SMoE 0.787 0.732 0.762 0.614 0.673 0.187 0.430
HC-SMoE 0.604 0.530 0.500 0.399 0.465 0.077 0.271
Pruning Frequency 0.823 0.756 0.751 0.595 0.676 0.223 0.449
EAN 0.805 0.744 0.754 0.601 0.672 0.209 0.441
REAP 0.841 0.768 0.762 0.624 0.696 0.248 0.472
GLM-4.5-Air Baseline 0.848 0.829 0.860 0.743 0.786 0.374 0.580
25% Merging M-SMoE 0.866 0.793 0.807 0.659 0.726 0.330 0.528
HC-SMoE 0.872 0.805 0.825 0.669 0.737 0.363 0.550
Pruning Frequency 0.848 0.811 0.854 0.706 0.759 0.341 0.550
EAN 0.872 0.817 0.876 0.720 0.768 0.374 0.571
REAP 0.884 0.829 0.839 0.688 0.759 0.412 0.585
50% Merging M-SMoE 0.518 0.500 0.519 0.437 0.468 0.099 0.284
HC-SMoE 0.707 0.659 0.706 0.577 0.618 0.220 0.419
Pruning Frequency 0.628 0.573 0.534 0.450 0.511 0.104 0.308
EAN 0.841 0.780 0.807 0.661 0.721 0.253 0.487
REAP 0.823 0.780 0.712 0.577 0.679 0.352 0.515
Qwen3-Coder- 480B-A35B- Instruct-FP8 Baseline 0.951 0.890 0.923 0.791 0.841 0.431 $\pm$ 0.011 0.636
25% Pruning Frequency 0.884 0.805 0.810 0.669 0.737 0.296 $\pm$ 0.017 0.516
EAN 0.939 0.878 0.911 0.775 0.827 0.419 $\pm$ 0.015 0.623
REAP 0.957 0.890 0.917 0.772 0.831 0.416 $\pm$ 0.013 0.624
50% Pruning Frequency 0.020 0.012 0.007 0.003 0.007 0.012 $\pm$ 0.001 0.010
EAN 0.915 0.841 0.854 0.714 0.777 0.382 $\pm$ 0.012 0.580
REAP 0.939 0.872 0.910 0.772 0.822 0.415 $\pm$ 0.015 0.619
Kimi-K2- Instruct- W4A16 Baseline 0.963 0.921 0.913 0.735 0.828 0.434 0.631
25% Pruning Frequency 0.530 0.463 0.595 0.508 0.486 0.082 0.284
EAN 0.909 0.860 0.857 0.698 0.779 0.379 0.579
REAP 0.957 0.921 0.918 0.759 0.840 0.440 0.640
50% Pruning Frequency 0.098 0.079 0.175 0.146 0.112 0.000 0.056
EAN 0.866 0.811 0.780 0.632 0.722 0.253 0.487
REAP 0.915 0.884 0.899 0.754 0.819 0.429 0.624
@@
ERNIE-4.5-21B- A3B-PT Baseline 0.815 0.231 0.523 0.564 0.782 0.873 0.813 0.737 0.462 0.812 0.724 0.721
25% Merging M-SMoE 0.061 0.016 0.039 0.497 0.729 0.860 0.723 0.602 0.424 0.801 0.699 0.667
HC-SMoE 0.369 0.099 0.234 0.515 0.728 0.860 0.745 0.649 0.428 0.794 0.694 0.677
Pruning Frequency 0.254 0.000 0.127 0.515 0.735 0.841 0.719 0.588 0.382 0.791 0.683 0.657
EAN 0.262 0.000 0.131 0.528 0.750 0.853 0.790 0.558 0.442 0.783 0.706 0.676
REAP 0.212 0.055 0.133 0.553 0.784 0.843 0.775 0.635 0.454 0.798 0.708 0.694
@@
REAP 0.015 0.000 0.008 0.403 0.562 0.713 0.668 0.391 0.388 0.708 0.669 0.563
Qwen3-30B-A3B Baseline 0.814 0.302 0.558 0.563 0.790 0.887 0.778 0.779 0.454 0.816 0.702 0.721
@@
REAP 0.329 0.104 0.217 0.404 0.616 0.828 0.643 0.517 0.360 0.632 0.637 0.580
Mixtral-8x7B- Instruct-v0.1 Baseline 0.469 0.123 0.296 0.650 0.842 0.887 0.861 0.691 0.496 0.722 0.740 0.736
25% Merging M-SMoE 0.296 0.044 0.170 0.532 0.775 0.828 0.746 0.529 0.424 0.603 0.632 0.634
HC-SMoE 0.385 0.121 0.253 0.608 0.811 0.876 0.838 0.631 0.484 0.736 0.726 0.714
Pruning Frequency 0.370 0.070 0.220 0.612 0.816 0.868 0.836 0.593 0.482 0.675 0.739 0.703
EAN 0.360 0.092 0.226 0.613 0.814 0.875 0.842 0.613 0.498 0.690 0.733 0.710
REAP 0.371 0.088 0.230 0.590 0.810 0.878 0.835 0.638 0.468 0.736 0.710 0.708
@@
HC-SMoE 0.152 0.033 0.093 0.540 0.764 0.862 0.795 0.544 0.448 0.675 0.709 0.667
Pruning Frequency 0.156 0.008 0.082 0.504 0.739 0.793 0.771 0.463 0.426 0.675 0.646 0.627
EAN 0.127 0.008 0.068 0.550 0.756 0.842 0.804 0.529 0.460 0.726 0.716 0.673
REAP 0.121 0.022 0.071 0.531 0.779 0.869 0.787 0.543 0.460 0.773 0.697 0.680
@@
Table A10: Additional $\tau^{2}$-bench results (passˆ1 scores) across non-reasoning and reasoning models and compression levels with REAP. Interleaved thinking is not applied for MiniMax-M2 (thinking traces from previous turns are discarded in multi-turn trajectories).
Model Dataset Baseline 20% 25% 30% 40%
Qwen3-Coder-30B-A3B Airline 0.393 0.407
Retail 0.626 0.620
Telecom 0.336 0.322
GLM-4.5-Air Airline (Thinking off) 0.633 0.640
Retail (Thinking off) 0.728 0.751
Telecom (Thinking off) 0.284 0.307
Telecom (Thinking on) 0.272 0.269
MiniMax-M2 Telecom 0.591 0.576 0.591 0.553
Table A11: Additional BFCLv3 results across across non-reasoning and reasoning models and compression levels with REAP. Interleaved thinking is not applied for MiniMax-M2 (thinking traces from previous turns are discarded in multi-turn trajectories).
Model Benchmark Baseline 20% 25% 30% 40%
Qwen3-Coder-30B-A3B BFCLv3 0.632 0.622
GLM-4.5-Air BFCLv3 (Thinking) 0.768 0.763
GLM-4.6-FP8 BFCLv3 (Thinking) 0.784 0.773 0.768 0.742
MiniMax-M2 BFCLv3 0.626 0.615 0.599 0.579
Table A12: Kimi-Linear-48B-A3B-Instruct results at 30% REAP compression across coding, math, and long-context evaluations.
Coding Math Long-Context
Model Compression HumanEval+ MBPP+ LiveCodeBench MATH-500 GSM8K LongBench v2 FRAMES
Kimi-Linear-48B-A3B-Instruct Baseline 0.823 0.669 0.276 0.818 0.873 0.368 0.557
30% 0.811 0.693 0.302 0.808 0.858 0.372 0.523
Figure˜A6 plots non-agentic coding and MC accuracy versus compressed model size. Figure˜7(a) depicts the proportion of singleton clusters for HC-SMoE and M-SMoE. Figure˜7(b) plots accuracy vs. maximum cluster sizes when the maximum cardinality of clusters is restricted. Figures˜A8 and A9 show the importance of using domain-specific calibration data, particularly at high compression ratios.
Table˜A9 presents the complete $\tau^{2}$-bench results across three domains (Retail, Airline, and Telecom) for the baseline model and REAP compression at 25% and 50% levels. The results show passˆk metrics for k=1, 2, and 3, demonstrating the impact of pruning on evaluating conversational agents, specifically designed to test their ability to collaborate with a user in real-world scenarios.
Table˜A10 presents additional $\tau^{2}$-bench results with REAP compression for a set of non-reasoning and reasoning models (Qwen3-Coder-30B, GLM4.5-Air, MiniMax-M2). Furthermore, Table˜A11 presents BFCLv3 scores for the same set of REAP-compressed models (with an inclusion of REAP-compressed GLM-4.6-FP8 variants in thinking mode). The calibration data mix used for non-reasoning models in these experiments is the same as for the other large-scale models in this study (12,228 samples including tool calling and agentic trajectory data). For the models that support reasoning, we include an extra 12,228 samples from the Open-R1 Mixture-of-Thoughts dataset (HuggingFace, 2025), drawing 4,096 samples from the Code, Math and Science subsets with a maximum sequence length of 16,384 tokens.
Table˜A12 shows results for the Kimi-Linear-48B-A3B-Instruct model as an example of REAP compression applied to a hybrid linear attention architecture. The calibration data mix is made up of 24,414 samples with a maximum sequence length of 16,384 tokens. The samples are drawn from evol-codealpaca (Chaudhary, 2023; Luo et al., 2024; Tam, 2023), xlam-function-calling (Liu et al., 2024c; Salesforce, 2025), SWE-smith-trajectories (Yang et al., 2025c; b), WildChat (Zhao et al., 2024), NuminaMath (LI et al., 2024) and Aya (Singh et al., 2024). The calibration data mix composition reflects coding, tool calling, multilingual and general conversational use cases. We evaluate Kimi-Linear-48B-A3B-Instruct on coding and math benchmarks at 30% REAP compression, as well as on long-context question-answering benchmarks LongBench v2 (Bai et al., 2025) and FRAMES (Krishna et al., 2024). For long-context benchmarks, a truncation in the middle is done for input prompts exceeding 131,072 tokens in length. The accuracy scores for FRAMES are estimated with LLM-as-a-judge (gpt-4.1-2025-04-14 model as judge). Notably, even though the maximum sequence length during calibration is 16,384 tokens, the REAP-compressed model retains performance at context lengths far exceeding that value.
@@
Figure A7: (7(a)) Average proportion of singleton clusters vs. model for HC-SMoE and M-SMoE. We find that the clustering algorithms used by our baseline merging methods tend to generate a high proportion of singleton clusters containing just a single expert. In order to achieve the desired compression ratio, the large number of singletons conversely results in some clusters which contain many experts, in some cases $N/2+1$ experts for a layer with $N$ experts are grouped into a single cluster. (7(b)) Accuracy vs. maximum cluster size using M-SMoE to compress 50% of experts in Qwen3-30B. While MC accuracy remains stable up to a maximum cluster size of 4, generative coding capabilities are severely diminished by restricting the clustering algorithm.